OpenStudy (anonymous):

@baby_bear69 can u help me again

OpenStudy (baby_bear69):

yep!

OpenStudy (anonymous):

The sum of d and 67 is equal to f. Let f = 83. Which equation can be used to find the value of d? A. d = 83 + 67 B. d + 67 = 83 C. d – 83 = 67 D. d – 67 = 83

OpenStudy (cggurumanjunath):

B

OpenStudy (baby_bear69):

well, start by filling in the letters for each equation

OpenStudy (anonymous):

can u help me with 3 more u guys please

OpenStudy (anonymous):

i think its b

OpenStudy (anonymous):

is b right

OpenStudy (baby_bear69):

you are correct(:

OpenStudy (anonymous):

can u help me with 3 more

OpenStudy (baby_bear69):

yes!

OpenStudy (anonymous):

The sum of j and m is equal to 58. Let j = 19. What is the value of m?

OpenStudy (anonymous):

subtract 19 from 58 :)

OpenStudy (anonymous):

u cant do taht it will be a negative

OpenStudy (anonymous):

no I mean take away 19 from 58 so 58-19

OpenStudy (anonymous):

its 39 i got it thx

OpenStudy (anonymous):

can u help me with 3 more or no ?

OpenStudy (anonymous):

:) awesome job!!!

OpenStudy (anonymous):

ummm sure if you'll fan and medal for me :D

OpenStudy (anonymous):

babybear u letf me

OpenStudy (baby_bear69):

post another question and tag me

OpenStudy (anonymous):

@helpblahblahblah look my ss is 15 XD XD XD apparently my fb account made me another os account without my knowledge :P